# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ 6 Conclusion
Motivated by the human-AI interplay in online content moderation, we propose a learning to defer model with limited and time-varying human capacity. In particular, for each period, the AI makes three decisions: (i) whether to keep or remove a new post (classification); (ii) whether to admit this new post for human review (admission); and (iii) which post to send to the next available reviewer (scheduling). The cost of a post is unknown until a human reviews it. The objective is to minimize the total loss with respect to an omniscient benchmark that knows the cost of posts. Since achieving vanishing loss is unattainable, we aim to minimize the average regret of a policy, i.e., the difference between the loss of a policy and a fluid benchmark. When the average cost of posts is known, we propose BACID which balances the idiosyncrasy loss avoided by admitting a post and the delay loss the admission could incur to other posts. We show that BACID achieves a near-optimal $O(\sqrt{\ell_{\max}})$ regret, with $\ell_{\max}$ being the maximum lifetime of a post. When the average cost of posts is unknown, we show that an optimism-only extension of BACID fails to learn because of the selective sampling nature of the system. That is, humans only see posts that are admitted by the AI, while labels from humans affect AI’s classification and admission accuracy. To address this issue, we carefully balance label-driven admissions and admissions aiming to reduce idiosyncrasy loss. Finally, we extend our algorithm to a contextual setting an derive a performance guarantee without dependence on the number of types. Our type aggregation technique for a many-class queueing system and analysis for queueing-delayed feedback enable us to provide (to the best of our knowledge) the first online learning result in contextual queueing systems, which may be of independent interest.
Our work also opens up a list of interesting questions. First, we assume reviewers produce perfect labels; how can we schedule posts when reviewers have skill-dependent non-perfect review quality? Second, our algorithm requires the knowledge of post lifetime, views per-period, and assumes constant views per-period. Can we relax these assumptions? Third, we focus on the $w$-fluid benchmark to handle non-stationary arrivals and human capacities, which is not necessary the tightest benchmark. Is there an alternative benchmark to consider? Fourth, we assume stationary expected cost of posts for the learning problem. It is interesting to extend our work to a non-stationary learning setting. Finally, our work focuses on the human-AI interplay in online content moderation and it is interesting to study this interplay in settings beyond content moderation.
@@ Appendix A Further discussion on content moderation practice (Section 1)
Although [MSA+21, ABB+22] do not provide a formal model, their views of the system are largely captured by our model. In particular, the severity of a content is modelled by the cost $c_{j}$ of a post in our model. Their admission policy, which removes posts that ML views as unambiguously harmful, and admits posts with predicted positive severity, can be viewed as a static two-threshold admission policy. Delaying reviews of harmful contents is undesirable for the system, as emphasized by our objective and the IV metric in [ABB+22]. We note that our paper does not aim to capture the full picture of content moderation. We assume that average cost is unknown but stationary, that per-period views of posts do not change over time, that we know the per-period views, the lifetime, and the idiosyncratic variance and that human reviewers are perfect (see Section 6 for further discussion). Approaches in [MSA+21, ABB+22] do not rely on these assumptions. However, the goal of our paper is to provide a stylized model that illustrates challenges arising from idiosyncratic variance in posts, time-varying capacity and selective sampling, and that relying on sample average and optimism-only learning policies may be inefficient in this setting. These justify the contribution of our paper which presents a formal model to address the two challenges.
@@ Appendix B Supplementary materials on BACID (Section 3)
B.2 Formulae to calculate per-period loss (Eq. (5))
Proposition 4.
For any normal random variable $X$ with mean $h$ and variance $\sigma^{2}$, we have
$\mathbb{E}\left[X^{+}\right]=h\Phi(h/\sigma)+\sigma\varphi(h/\sigma)\qquad\text{and}\qquad\mathbb{E}\left[X^{-}\right]=-h\Phi(-h/\sigma)+\sigma\varphi(-h/\sigma).$
@@
We prove the result for $\mathbb{E}\left[X^{-}\right]$. The result for $\mathbb{E}\left[X^{+}\right]$ follows by considering $\mathbb{E}\left[(-X)^{-}\right]$. We have $\mathbb{E}\left[X^{-}\right]=-\mathbb{P}\{X\leq 0\}\mathbb{E}\left[X\,\middle|\,X\leq 0\right]$. Conditioning on $X\leq 0$, $X$ has a truncated normal distribution; the mean $\mathbb{E}\left[X\,\middle|\,X\leq 0\right]$ is thus $h+\frac{-\varphi(-h/\sigma)}{\Phi(-h/\sigma)}\sigma$. Since $\mathbb{P}\{X\leq 0\}=\Phi(-h/\sigma)$, we have
$\mathbb{E}\left[X^{-}\right]=-\Phi(-h/\sigma)\left(h+\frac{-\varphi(-h/\sigma)}{\Phi(-h/\sigma)}\sigma\right)=-h\Phi(-h/\sigma)+\sigma\varphi(-h/\sigma).$
B.3 Proof of the Lower Bound (Theorem 2)
Proof.
Consider a setting where there is only one type of posts ($K=1$). Since there is only one post, we will omit notational dependence on the type. The arrival rate is $\lambda=1/2$ and the service rate is $\mu=1/2$. We assume $N(t)=1$. For this type, $h=0$ and $\sigma=1$; so the inaccuracy loss per post is $r=1$ by Lemma 4. The lifetime is $\ell$. Recall $D(j)$ is the delay for a post to review. Let $Q(t)$ be the length of the review queue at the beginning of period $t$. Then for a policy $\pi$,
@@ B.3 Proof of the Lower Bound (Theorem 2)
B.4 Bounding idiosyncrasy loss by Lagrangian via drift analysis (Lemma 3.3)
@@
B.5 Connecting Lagrangian with Primal Objective (Lemma 3.5)
@@
B.6 Difference in Lagrangian by Having Window-Based Duals (Lemma 3.6)
@@
B.7 BACID optimizes per-period Lagrangian (Lemma 3.7)
@@ Appendix C Supplementary materials on OLBACID (Section 4)
C.1 Solution to the optimistic estimator (Footnote 4)
Lemma C.1.
For any $h_{1}<h_{2}$ and $k\in\mathcal{K}$, the solution $h^{\star}$ of $\arg\max_{h\in[h_{1},h_{2}]}r_{k}(h)$ is as follows: if $0\in[h_{1},h_{2}]$, then $h^{\star}=0$; if $0<h_{1}$, then $h^{\star}=h_{1}$; if $h_{2}<0$, then $h^{\star}=h_{2}.$
Proof.
By the definition of $r_{k}^{O}$ and $r_{k}^{R}$ in (5), we know for $h>0$ that
$r_{k}^{R}(h)=v_{k}(-h\Phi(-h/\sigma_{k})+\sigma_{k}\varphi(-h/\sigma_{k}))$ $=v_{k}(-h\Phi(-h/\sigma_{k})+\sigma_{k}\varphi(h/\sigma_{k}))$
$\leq v_{k}(h\Phi(h/\sigma_{k})+\sigma_{k}\varphi(h/\sigma_{k}))=r_{k}^{O}(h).$
Similarly, we have $r_{k}^{O}(h)\leq r_{k}^{R}(h)$ for $h<0$. Therefore, $r_{k}(h)=r_{k}^{O}(h)$ for $h<0$, $r_{k}(h)=r_{k}^{R}(h)$ for $h>0$, and $r_{k}(0)=r_{k}^{O}(0)=r_{k}^{R}(0)$. In addition, $r_{k}^{R}(h)$ is decreasing for $h\geq 0$ and $r_{k}^{O}(h)$ is increasing for $h\leq 0$ by Fact 7. As a result, $r_{k}(0)\geq r_{k}(h)$ for any $h$, so $h^{\star}=0$ if $0\in[h_{1},h_{2}]$. The remaining result follows from the monotonicity properties of $r_{k}^{R}$ and $r_{k}^{O}$. ∎
C.2 Optimism-only fails to learn classification decisions (Proposition 3)
We recall the setting: $K=2$, $h_{1}=0,h_{2}=-1,v_{1}=\sigma_{1}=v_{2}=\sigma_{2}=1$ and $\ell_{1}=49\ell_{2}$. The algorithm is run with $H=1$, with exact knowledge that $h_{1}=0$, and with no knowledge of $h_{2}$. The hyper-parameter $\beta$ is in $(48/\ell_{1},1)$. The time horizon $T$ is $\lceil\exp\left(\beta\ell_{1}/576\right)\rceil$. Arrival rates are such that $\lambda_{1}(t)=1,\lambda_{2}(t)=0$ for $t\leq T_{1}\coloneqq\lfloor\frac{\beta\ell_{1}}{2}\rfloor$, and $\lambda_{1}(t)=5/6,\lambda_{2}(t)=1/6$ for $T_{1}<t\leq T$. The service rate is $\mu_{1}=\mu_{2}=1/2$ with $N(t)=1$ for any $t$. We know that $r_{2}^{R}-r_{2}^{O}=\Phi(1)+\varphi(1)-(\Phi(-1)+\varphi(-1))=\Phi(1)-\Phi(-1)>0$, meaning that wrongly removal of a type-$2$ post incurs an additional loss of $(\Phi(1)-\Phi(-1))\ell_{2}$. Let us define $Q_{\max}=\beta\ell_{1}$.
@@ C.2 Optimism-only fails to learn classification decisions (Proposition 3)
We have $r_{1}=\min(r_{1}^{O},r_{1}^{R})=1,r_{2}=\min(r_{2}^{O},r_{2}^{R})\leq 1$ by (5). In addition, the admission rule is that $A_{k}(t)=1$ if $\beta\bar{r}_{k}(t)\ell_{k}>Q_{k}(t)$. Since $\bar{r}_{k}(t)\leq\sigma_{k}=1$ by Fact 7, it holds that $A_{2}(t)\leq\mathbbm{1}\left(Q_{2}(t)\leq\beta\ell_{2}\right)$, which, by the same analysis in the proof of Lemma 3.2, implies that $Q_{2}(t)\leq\beta\ell_{2}+1<\frac{\beta\ell_{1}}{24}=\frac{Q_{\max}}{24}$ because $\ell_{1}\geq 49\ell_{2}$ and $\beta\geq\frac{48}{\ell_{1}}.$
@@
We fix a period $t>T_{1}$ and look at the interval $[t-E,t-1]$. Since $Q_{\max}=\beta\ell_{1}\geq 24$, we have $E<Q_{\max}/2$. We consider two cases:
@@
C.3 Number of posts admitted to label-driven queue (Lemma 4.4)
@@
C.4 Regret guarantee for OLBACID (Theorem 3)
@@
We thus focus on $T\geq 3$, $K\leq\ell_{\max}$ and $\beta=\gamma=1/\sqrt{K\ell_{\max}}\geq 1/\ell_{\max}$, which is the setting Lemmas 4.14.2 and 4.3 are applicable. Omitting dependence on $v_{\max},r_{\max},\sigma_{\max},\hat{\mu}_{\min},\ln T$, constants and applying Lemmas 4.14.2 and 4.3 to the corresponding terms in (12), we obtain:
@@
C.5 Confidence intervals are valid with high probability (Lemma 4.5)
@@
Fix $k\in\mathcal{K},t\in[T]$. The number $n_{k}(t)$ of type-$k$ posts in the dataset $\mathcal{D}_{t}$ is upper bounded by $t$. Conditioning on $n_{k}(t)=n\leq t$, there are $n$ posts whose costs are i.i.d. normal with mean $h_{k}$ and variance $\sigma_{k}^{2}$. By Chernoff Bound (Fact 2), the empirical average $\hat{h}_{k}(t)$ satisfies
$\mathbb{P}\left\{|\hat{h}_{k}(t)-h_{k}|>\sigma_{k}\sqrt{\frac{8\ln t}{n}}\right\}\leq 2\exp\left(\frac{-8\sigma_{k}^{2}\ln t\cdot n}{2\sigma_{k}^{2}\cdot n}\right)\leq\frac{2}{t^{4}}.$
By union bound over $n=0,\ldots,t-1$, this bound is simultaneously valid for all values of $n_{k}(t)$, and thus $\mathbb{P}\left\{|\hat{h}_{k}(t)-h_{k}|>\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right\}\leq 2t^{-3}$. We finish the proof by using the definition of the confidence interval in (11) and the assumption that $h_{k}\in[-H,H]$. ∎
C.6 Bounding number of admitted posts to label-driven queue (Lemma 4.6)
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)(\mathbbm{1}\left(\mathcal{E}_{k,t}+\mathcal{E}_{k,t}^{c}\right))\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+\sum_{t=1}^{T}\frac{2}{t^{3}}\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+4$
where the first inequality uses Lemma 4.5 and the fact that $\sum_{k\in\mathcal{K}}E_{k}(t)\leq 1$, and the second inequality uses $\sum_{t=1}^{\infty}1/t^{3}\leq 2$. It remains to bound the first sum.
Fixing a type $k$, we consider the last time period $T_{k}$ such that $E_{k}(T_{k})\mathbbm{1}\left(\mathcal{E}_{k,T_{k}}\right)=1$. By Line 2 in Algorithm 2, $E_{k}(T_{k})=1$ if and only if $\underline{h}_{k}(T_{k})<-\gamma<\gamma<\bar{h}_{k}(T_{k})$ and $Q^{\mathcal{\textsc{LD}}}(T_{k})=0$. In addition, the setting of the confidence interval, (11), shows that $\bar{h}_{k}(T_{k})-\underline{h}_{k}(T_{k})\leq 2\sigma_{k}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$ where $n_{k}(T_{k})$ is the number of type-$k$ posts in the dataset $\mathcal{D}(T_{k})$. We next bound $n_{k}(T_{k})$ for the case of $h_{k}\geq 0$. Since $\mathbbm{1}\left(\mathcal{E}_{k,T_{k}}\right)=1$, we have $h_{k}\in[\underline{h}_{k}(T_{k}),\bar{h}_{k}(T_{k})]$. by definition,
$\underline{h}_{k}(T_{k})=\bar{h}_{k}(T_{k})-(\bar{h}_{k}(T_{k})-\underline{h}_{k}(T_{k}))\geq h_{k}-2\sigma_{k}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}\geq\eta-2\sigma_{k}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$
where the last inequality is by the margin definition that $\eta\leq|h_{k}|=h_{k}$. Since $\underline{h}_{k}(T_{k})<-\gamma$, we have $-\gamma\geq\eta-2\sigma_{k}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$. As a result, $n_{k}(T_{k})\leq\frac{32\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$. The same analysis applies for $h_{k}<0$.

### v3

@@ 7 Conclusion
Motivated by the human-AI interplay in online content moderation, we propose a learning to defer model with limited and time-varying human capacity. For each period, the AI makes three decisions: (i) keep or remove a new post (classification); (ii) admit this new post for human review (admission); and (iii) select a post for the next available reviewer (scheduling). The cost of a post is unknown until a human reviews it. The objective is to minimize the total loss with respect to an omniscient benchmark that knows the cost of posts. Since achieving vanishing loss is unattainable, we aim to minimize the average regret of a policy, i.e., the difference between the loss of a policy and a fluid benchmark. When the cost distribution of posts is known, we propose BACID which balances the idiosyncrasy loss avoided by admitting a post and the delay loss the admission could incur to other posts. We show that BACID achieves a near-optimal $O(\sqrt{\ell_{\max}})$ regret, with $\ell_{\max}$ being the maximum lifetime of a post. When the cost distribution of posts is unknown, we show that an optimism-only extension of BACID fails to learn because of the selective sampling nature of the system. That is, humans only see posts admitted by the AI, while labels from humans affect AI’s classification accuracy. To address this issue, we introduce label-driven admissions. Finally, we extend our algorithm to a contextual setting and derive a regret guarantee without dependence on the number of types. Our type aggregation technique for a many-class queueing system and analysis for queueing-delayed feedback enable us to provide (to the best of our knowledge) the first online learning result in contextual queueing systems, which may be of independent interest.
Our work also opens up a list of interesting questions. First, we assume reviewers produce perfect labels; how can we schedule posts when reviewers have skill-dependent non-perfect review quality? Second, our algorithm requires the knowledge of post lifetime and assumes constant per-period cost throughout a post’s lifetime. Can we relax these assumptions? Third, we focus on the $w$-fluid benchmark to handle non-stationary arrivals and human capacities, which is not necessary the tightest benchmark. Is there an alternative benchmark to consider? Fourth, although we allow for time-varying arrival patterns and human reviewing capacity, we assume stationary cost distributions of posts for the learning problem. It is interesting to extend our work to a non-stationary learning setting. Finally, our work focuses on the human-AI interplay in content moderation and it is interesting to study this interplay in settings beyond content moderation.
@@ Appendix A Further discussion on content moderation practice (Section 1)
Although [MSA+21, ABB+22] do not provide a formal model, their views of the system are largely captured by our model. In particular, the severity of a content is modelled by the cost $c_{j}$ of a post in our model. Their admission policy, which removes posts that ML views as unambiguously harmful, and admits posts with predicted positive severity, can be viewed as a static two-threshold admission policy. Delaying reviews of harmful contents is undesirable for the system, as emphasized by our objective and the IV metric in [ABB+22]. We note that our paper does not aim to capture the full picture of content moderation. We assume that cost distributions are unknown but stationary, that per-period cost of a post does not change over time, that we know the lifetime and that human reviewers are perfect (see Section 7 for further discussion). Approaches in [MSA+21, ABB+22] do not rely on these assumptions. However, the goal of our paper is to provide a stylized model that illustrates challenges arising from idiosyncratic uncertainty in posts, time-varying capacity and selective sampling, and that relying on sample average and optimism-only learning policies may be inefficient in this setting. These justify the contribution of our paper which presents a formal model to address the two challenges.
@@ Appendix B Supplementary materials on BACID (Section 3)
B.2 Proof of the Lower Bound (Theorem 2)
@@
Consider a setting where there is only one type of posts ($K=1$) and, in any period $t$, we only have one reviewer, i.e., $N(t)=1$. Since there is only one post, we will omit notational dependence on the type. The arrival rate is $\lambda=1/2$ and the service rate is $\mu=1/2$. We assume $r=1$ and denote the lifetime by $\ell$. Recall that $D(j)$ is the delay for a post to review. Let $Q(t)$ be the length of the review queue at the beginning of period $t$. Then for a policy $\pi$,
@@
B.3 Bounding idiosyncrasy loss by Lagrangian via drift analysis (Lemma 3.3)
@@
B.4 Connecting Lagrangian with Primal Objective (Lemma 3.5)
@@
B.5 Difference in Lagrangian by Having Window-Based Duals (Lemma 3.6)
@@
B.6 BACID optimizes per-period Lagrangian (Lemma 3.7)
@@ Appendix C Supplementary materials on OLBACID (Section 4)
C.1 Optimism-only fails to learn classification decisions (Proposition 3)
We recall the setting: $K=2$, $\ell_{1}=96\ell_{2}$ and
$\mathcal{F}_{1}=\{\pm 1~{}\text{with probability 0.5}\},\mathcal{F}_{2}=\{-1~{}\text{with probability 0.05};~{}+1~{}\text{with probability 0.95}\}.$
The algorithm is run with $r_{\max}=1$, with exact knowledge of $\mathcal{F}_{1}$, and with no knowledge of $\mathcal{F}_{2}$. The hyper-parameter $\beta$ is in $(96/\ell_{1},1)$. The time horizon $T$ is $\lceil\exp\left(\beta\ell_{1}/576\right)\rceil$. Arrival rates are such that $\lambda_{1}(t)=1,\lambda_{2}(t)=0$ for $t\leq T_{1}\coloneqq\lfloor\frac{\beta\ell_{1}}{2}\rfloor$, and $\lambda_{1}(t)=5/6,\lambda_{2}(t)=1/6$ for $T_{1}<t\leq T$. The service rate is $\mu_{1}=\mu_{2}=1/2$ with $N(t)=1$ for any $t$. We know that $r_{2}^{O}-r_{2}^{R}=0.95-0.05=0.9$, meaning that wrongly keeping of a type-$2$ post incurs an additional loss of $0.9\ell_{2}$. Let us define $Q_{\max}=r_{1}\beta\ell_{1}=0.5\beta\ell_{1}$.
@@ C.1 Optimism-only fails to learn classification decisions (Proposition 3)
We have $r_{1}=\min(r_{1}^{O},r_{1}^{R})=0.5$. In addition, the admission rule is that $A_{k}(t)=1$ if $\beta\bar{r}_{k}(t)\ell_{k}>Q_{k}(t)$. Since $\bar{r}_{k}(t)\leq r_{\max}=1$, it holds that $A_{2}(t)\leq\mathbbm{1}\left(Q_{2}(t)\leq\beta\ell_{2}\right)$, which, by the same analysis in the proof of Lemma 3.2, implies that $Q_{2}(t)\leq\beta\ell_{2}+1<\frac{\beta\ell_{1}}{48}=\frac{Q_{\max}}{24}$ because $\ell_{1}\geq 96\ell_{2}$ and $\beta>\frac{96}{\ell_{1}}.$
@@
We fix a period $t>T_{1}$ and look at the interval $[t-E,t-1]$. Since $Q_{\max}=r_{1}\beta\ell_{1}\geq 24$, we have $E<Q_{\max}/2$. We consider two cases:
@@
C.2 Number of posts admitted to label-driven queue (Lemma 4.4)
@@
C.3 Regret guarantee for OLBACID (Theorem 3)
@@
We thus focus on $T\geq 3$, $K\leq\ell_{\max}$ and $\beta=\gamma=1/\sqrt{K\ell_{\max}}\geq 1/\ell_{\max}$, which is the setting Lemmas 4.14.2 and 4.3 are applicable. Omitting dependence on $r_{\max},\sigma_{\max},\hat{\mu}_{\min},\ln T$, constants and applying Lemmas 4.14.2 and 4.3 to the corresponding terms in (13), we obtain:
@@
C.4 Confidence bounds are valid with high probability (Lemma 4.5)
We first need the following result that a Lipschitz function of a sub-Gaussian random variable is still sub-Gaussian.
Lemma C.3.
Given a sub-Gaussian random variable $X$ with variance proxy $\sigma^{2}$ and a $1-$Lipschitz function $f(\cdot)$, the random variable $f(X)$ is sub-Gaussian with variance proxy $2\sigma^{2}.$
Proof.
The proof follows a simplified version of the proof of [Kon14, Theorem 1]. Define $V=f(X)-\mathbb{E}_{X^{\prime}}\left[f(X^{\prime})\right]$ where $X^{\prime}$ is an independent copy of $X$. By definition, showing that $f(X)$ is sub-Gaussian with variance proxy $2\sigma^{2}$ is equivalent to showing that $\mathbb{E}_{X}\left[\exp(sV)\right]\leq\exp(s^{2}(2\sigma^{2})/2)$ for any $s\in\mathbb{R}.$ Fixing $s\in\mathbb{R}$ and expanding the term gives
$\mathbb{E}_{X}\left[e^{sV}\right]$ $=\mathbb{E}_{X}\left[e^{s\mathbb{E}_{X^{\prime}}\left[f(X)-f(X^{\prime})\right]}\right]$
$\leq\mathbb{E}_{X,X^{\prime}}\left[e^{s(f(X)-f(X^{\prime}))}\right]$ (Jensen’s inequality)
$=\frac{1}{2}\mathbb{E}_{X,X^{\prime}}\left[e^{s(f(X)-f(X^{\prime}))}+e^{s(f(X^{\prime})-f(X))}\right]$
$\overset{(a)}{\leq}\frac{1}{2}\mathbb{E}_{X,X^{\prime}}\left[e^{s|X-X^{\prime}|}+e^{-s|X-X^{\prime}|}\right]$
=$\overset{(U=X-X^{\prime})}{=}\frac{1}{2}\mathbb{E}_{U}\left[e^{sU}+e^{-sU}\right]=\mathbb{E}_{U}\left[e^{sU}\right],$ (By the symmetry of $U$)
where inequality (a) uses the fact that $e^{a}+e^{-a}=2\mathrm{cosh}(a)\leq 2\mathrm{cosh}(b)=e^{b}+e^{-b}$ for any $b\geq|a|.$ Since $-X^{\prime}$ is sub-Gaussian with variance proxy $\sigma^{2}$ (Fact 6) and $X,X^{\prime}$ are independent, we have $U=X+(-X^{\prime})$ sub-Gaussian with variance proxy $2\sigma^{2}$ by Fact 7. As a result, $\mathbb{E}_{U}\left[e^{sU}\right]\leq\exp(s^{2}(2\sigma^{2})/2)$ and $\mathbb{E}_{X}\left[e^{sV}\right]\leq\mathbb{E}_{U}\left[e^{sU}\right]\leq\exp(s^{2}(2\sigma^{2})/2)$ for any $s\in\mathbb{R}$, showing that $f(X)$ is sub-Gaussian with variance proxy $2\sigma^{2}$. ∎
@@
Fix $k\in\mathcal{K},t\in[T]$. The number $n_{k}(t)$ of type-$k$ posts in the dataset $\mathcal{D}_{t}$ is upper bounded by $t$. Conditioning on $n_{k}(t)=n\leq t$, there are $n$ posts whose costs are i.i.d. samples from $\mathcal{F}_{k}.$ We first show that $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ with probability at least $1-2t^{-4}.$ Denote the observed costs by $X_{1},\ldots,X_{n}.$ By definition,
$\hat{h}_{k}(t)=\hat{r}_{k}^{O}(t)-\hat{r}_{k}^{R}(t)=\frac{\sum_{i=1}^{n}X_{i}^{+}-(\sum_{i=1}^{n}(-X_{i}^{-})}{n}=\frac{\sum_{i=1}^{n}X_{i}^{+}+\sum_{i=1}^{n}X_{i}^{-}}{n}=\frac{\sum_{i=1}^{n}X_{i}}{n}.$
Since $\mathcal{F}_{k}$ is sub-Gaussian with variance proxy $\sigma_{\max}^{2}$, applying Chernoff bound (Fact 2) gives
$\mathbb{P}\left\{|\hat{h}_{k}(t)-h_{k}|>\sigma_{\max}\sqrt{\frac{8\ln t}{n}}\,\middle|\,|n_{k}(t)=n\right\}\leq 2\exp\left(-4\ln t\right)=2t^{-4}.$
Combining it with the definitions of $\underline{h}_{k}(t)$ and $\bar{h}_{k}(t)$ in (11) then gives
$\mathbb{P}\left\{h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]\,\middle|\,|n_{k}(t)=n\right\}\geq 1-2t^{-4}.$ (21)
We next show $r_{k}\leq\bar{r}_{k}(t)$ with probability at least $1-2t^{-4}$ conditioned on $n_{k}(t)=n$. By definition, $\hat{r}_{k}^{O}(t)=\frac{\sum_{i=1}^{n}X_{i}^{+}}{n}.$ In addition, since $\{X_{i}\}_{i\leq n}^{\prime}$s are i.i.d. sub-Gaussian random variables with variance proxy $\sigma^{2}_{\max}$, we also have $\{X_{i}^{+}\}_{i\leq n}$’s are i.i.d. sub-Gaussian random variables with variance proxy $2\sigma^{2}_{\max}$ by Lemma C.3. Applying Chernoff bound (Fact 2) and the fact that $r_{k}^{O}=\mathbb{E}\left[X_{i}^{+}\right]$ gives
$\mathbb{P}\left\{r_{k}^{O}-\hat{r}_{k}^{O}(t)>4\sigma_{\max}\sqrt{\frac{\ln t}{n}}\,\middle|\,|n_{k}(t)=n\right\}\leq 2\exp\left(-4\ln t\right)=t^{-4}.$
By symmetry, we also have
$\mathbb{P}\left\{r_{k}^{R}-\hat{r}_{k}^{R}(t)>4\sigma_{\max}\sqrt{\frac{\ln t}{n}}\,\middle|\,|n_{k}(t)=n\right\}\leq 2\exp\left(-4\ln t\right)=t^{-4}.$
By union bound, the above two inequalities imply that
$\mathbb{P}\left\{\min(r_{k}^{R},r_{k}^{O})-\min(\hat{r}_{k}^{R}(t),\hat{r}_{k}^{O}(t))>4\sigma_{\max}\sqrt{\frac{\ln t}{n}}\,\middle|\,|n_{k}(t)=n\right\}\leq 2\exp\left(-4\ln t\right)=2t^{-4}$
Using the definition that $r_{k}=\min(r_{k}^{R},r_{k}^{O})$ and that of $\bar{r}_{k}(t)$ in (12) shows
$\mathbb{P}\{r_{k}\leq\bar{r}_{k}(t)|n_{k}(t)=n\}\geq 1-2t^{-4}.$
Combining it with (21) gives $\mathbb{P}\{\mathcal{E}_{k,t}|n_{k}(t)=n\}\geq 1-4t^{-4}$ by union bound. The lemma then follows by applying union bound over $n=0,\ldots,t-1.$
C.5 Bounding number of admitted posts to label-driven queue (Lemma 4.6)
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)(\mathbbm{1}\left(\mathcal{E}_{k,t}+\mathcal{E}_{k,t}^{c}\right))\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+\sum_{t=1}^{T}\frac{4}{t^{3}}\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+5$
where the first inequality uses Lemma 4.5 and the fact that $\sum_{k\in\mathcal{K}}E_{k}(t)\leq 1$, and the second inequality uses $\sum_{t=1}^{\infty}1/t^{3}\leq 1.2$. It remains to bound the first sum.
Fixing a type $k$, we consider the last time period $T_{k}$ such that $E_{k}(T_{k})\mathbbm{1}\left(\mathcal{E}_{k,T_{k}}\right)=1$. By Line 2 in Algorithm 2, $E_{k}(T_{k})=1$ if and only if $\underline{h}_{k}(T_{k})<-\gamma<\gamma<\bar{h}_{k}(T_{k})$ and $Q^{\mathcal{\textsc{LD}}}(T_{k})=0$. In addition, the setting of the confidence interval, (11), shows that $\bar{h}_{k}(T_{k})-\underline{h}_{k}(T_{k})\leq 2\sigma_{\max}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$ where $n_{k}(T_{k})$ is the number of type-$k$ posts in the dataset $\mathcal{D}(T_{k})$. We next bound $n_{k}(T_{k})$ for the case of $h_{k}\geq 0$. Since $\mathbbm{1}\left(\mathcal{E}_{k,T_{k}}\right)=1$, we have $h_{k}\in[\underline{h}_{k}(T_{k}),\bar{h}_{k}(T_{k})]$. by definition,
$\underline{h}_{k}(T_{k})=\bar{h}_{k}(T_{k})-(\bar{h}_{k}(T_{k})-\underline{h}_{k}(T_{k}))\geq h_{k}-2\sigma_{\max}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}\geq\eta-2\sigma_{\max}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$
where the last inequality is by the margin definition that $\eta\leq|h_{k}|=h_{k}$. Since $\underline{h}_{k}(T_{k})<-\gamma$, we have $-\gamma\geq\eta-2\sigma_{\max}\sqrt{\frac{8\ln T_{k}}{n_{k}(T_{k})}}$. As a result, $n_{k}(T_{k})\leq\frac{32\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$. The same analysis applies for $h_{k}<0$.